OpenStudy (anonymous):

Write the standard form of the line that passes through the given points. (7, -3) and (4, -8)

OpenStudy (amorfide):

y=mx+c is the standard equation of a straight line to get this you must work out the y intercept, and the gradient the gradient of a line passing through two coordinates is the difference in y ordinates divided by the difference in x coordinates labelling the first set of coordinates to be (x1, y1) and the second set (x2, y2) the gradient would be equal to (y2-y1)/(x2-x1) so m=(y2-y1)/(x2-x1) next to get the equation of a line you use the formula y-y1=m(x-x1) where y1 is the y ordinate of one of the two coordinate sets, which ever one you choose, you must use the same set for the x ordinate for example if i did y-(-3) i would have to use 7 for the x1 to be consistent can you work out the answer? I will confirm if you are correct, or help along the way

OpenStudy (anonymous):

???

OpenStudy (amorfide):

labelling the first set of coordinates to be (x1, y1) and the second set (x2, y2) (x1,y1)= (7, -3) (x2,y2)= (4, -8) now work it out

OpenStudy (anonymous):

And that will be in standard form?

OpenStudy (amorfide):

you have to work it out, i am not going to give you the answer

OpenStudy (anonymous):

I'll get it, don't worry..lol